# A Linear Quantum Coupler for Clean Bosonic Control

### v2

@@ Appendix D Effects of parasitic loop inductance
where $l_{m}$, $s_{m}$ and $r_{m}$ are the m-th order derivatives of the potential energy functions of the left, shunting and right LINC branches, respectively, evaluated at ±$\pm\frac{\phi}{2}$. Finally, for driven behavior, we can also compute the n-th order derivative with respect to external flux, giving the general expansion coefficient
@@
{$U_{L,R}&=\frac{E_{\text{loop}}}{2}\left(\beta_{l}\sin\varphi_{J_{L,R}}\right)^{2}-E_{J}\cos\varphi_{J_{L,R}}\\ \phi_{L,R}&=\varphi_{J_{L,R}}+\beta_{l}\sin\varphi_{J_{L,R}}\end{aligned}\right.$ (50)
@@
It is then possible to graphically evaluate the exact expressions of the LINC Hamiltonian expansion coefficients. We perform this numerically to evaluate effects on the static LINC in LABEL:appx_fig4b, c. We expect the loop inductance of the LINC to be geometrically limited to $\sim 1$ pH/ $\mu$m, which for a $100\mu$m loop side implies $L_{\text{loop}}=0.2$nH $\implies\beta_{l}\sim 0.02$ (with $L_{S}\sim 3$nH, $L_{J}\sim 10$nH). We observe that the static frequency of the LINC as a function of flux is barely affected for in this range of loop inductance. This also implies that the parametric mixing strength $g_{3wm}\sim 0.5\;d\omega_{L}/d\phi_{d}$ remains similar. The static Kerr of the LINC is more significantly affected, with $\alpha_{L}=$1-10 MHz at $\phi_{DC}=\pi/2$, but the Kerr still has a zero-crossing very close to half flux, and can therefore always be nulled out for sensitive operations.
Appendix E Effect of a series parasitic inductance
We now consider the effect of a parasitic inductance in series with the LINC loop, which may arise from the geometric inductance of the metallic arms of the coupler that form the electric dipole necessary for self-capacitance and coupling. As in appendix A, we label the phase across the LINC as $\theta_{c}$ (with the flux threading it as $\phi_{d}$), with potential energy $U_{\mathrm{LINC}}$. We then consider the phase across the whole effective dipole (LINC $+$ parasitic inductor) as $\theta$, with corresponding potential energy $U_{\text{tot}}$. Each nonlinear process activated by the coupler is then described through derivatives of this total potential energy, and can be analytically derived. We will see that for a symmetric LINC, this parasitic inductance will primarily deform the potential and dilute the nonlinearity, but not shift the potential minima – thus preserving the LINC’s linearity at $\phi_{\text{DC}}=\pi/2$. We outline our derivation in a pedagogical manner below, with the actual analytic results for the LINC Hamiltonian detailed in section E.3.
@@ Appendix E Effect of a series parasitic inductance
To evaluate the correspondence in Eq. (52), we appeal to the current conservation condition $i=i_{\mathrm{LINC}}$ to relate first-order partial derivatives of $U$ and $U_{\mathrm{LINC}}$, following methodology similar to [29, 32, 65]. As $i=\frac{2\pi}{\Phi_{0}}\frac{\partial U}{\partial\theta}$, and $i_{\mathrm{LINC}}=\frac{2\pi}{\Phi_{0}}\frac{\partial U_{\mathrm{LINC}}}{\partial\theta_{c}}$, the following relation always holds:
@@
E.1 Representing the total potential energy and phase
In this section, we analyze the change of variables from the LINC variables $(\theta_{c},\phi_{d})$ to the effective dipole variables $(\theta_{c},\phi_{d})$ and study its invertibility. We define the bi-variate function implementing the change of variables as $\mathbf{F}$, such that $(\theta,\phi_{d})=\mathbf{F}(\theta_{c},\phi_{d})$. Let us define the superconducting phase drop across the parasitic inductor to be $\theta_{p}$, satisfying $\theta=\theta_{c}+\theta_{p}$, and its current to be $i_{p}$. Our derivation will primarily use the current conservation relation $i_{\mathrm{LINC}}=i_{p}$.
@@ E.1 Representing the total potential energy and phase
Applying current conservation and phase addition to eqs. 54 and 55, the effective dipole phase reads,
@@
E.2 Computing derivatives of the total potential
@@
[…]
@@
[…]
@@
$U_{mn}:=\frac{\partial^{m+n}U_{\text{tot}}}{\partial\theta_{c}^{m}\phi_{d}^{n}}$ (67)
@@
$U_{m+1n}=\frac{\partial^{m+n}}{\partial\theta_{c}^{m}\phi_{d}^{n}}(E_{L}p_{00}+E_{J}u_{10})$ (68)
@@
In combination with the expression for $p_{20}$ previously obtained, this gives an expression of $U_{12}$ that again only containes $p_{10}$ and $u_{mn}$ terms. Thus by recursively substituting the expressions for $p_{01}$, $p_{11}$ and $p_{20}$, any partial derivative of $U$ can be expressed as a multinomial function of $p_{10}$ and the set of nonlinear coefficients $u_{mn}$.
@@
E.3 The LINC Hamiltonian with parasitic inductance
Given the derivatives of the total potential $U_{\text{tot}}$ with respect to $\hat{\theta}$ and $\phi_{d}$, we can find the relevant coupler Hamiltonian in the presence of linear inductance, which we derive here. Similar to eq. 21, the Hamiltonian we derive describes the equations of motion corresponding to the bosonic operator $\hat{\theta}=\theta_{zpf}\left[\phi_{\text{DC}}\right](\hat{c}+\hat{c}^{\dagger})$, which can be expanded in terms of $U_{mn}$ as:
$\begin{split}&H_{\text{tot}}=\sum_{m,n}\tilde{g}_{mn}[\phi_{\text{DC}}]\;\phi^{n}_{AC}\;\left(\hat{c}+\hat{c}^{\dagger}\right)^{m}\\ &\tilde{g}_{mn}:=\frac{1}{m!\;n!}\left.\left(U_{m,n}\theta_{zpf}^{m}+U_{m,0}\;\frac{\partial^{n}\theta_{zpf}^{m}}{\partial\phi_{d}^{n}}\right)\right|_{\begin{subarray}{c}\hat{\theta}=0\\ \phi_{d}=\phi_{\text{DC}}\end{subarray}}\\ &\text{where }\theta_{zpf}=\left(\dfrac{2E_{C}}{U_{2,0}}\right)^{1/4}\end{split}$ (79)
Note that the parasitic inductance also affects $U_{2,0}$ and hence the spread of the wavefunction $\theta_{zpf}$, whose intrinsic flux dependence affects the every flux driven process.
First, let us consider the static Hamiltonian, given by:
$\begin{split}H_{\text{LINC}}&=\sum_{m}\tilde{g}_{m,0}\left(\hat{c}+\hat{c}^{\dagger}\right)^{m}\\ \tilde{g}_{m,0}&=U_{m,0}\left.\frac{\theta_{zpf}}{m!}\right|_{\phi_{\text{DC}}}.\end{split}$
@@ E.3 The LINC Hamiltonian with parasitic inductance
$\begin{gathered}U_{2,0}=p_{10}\left(E_{L}+2E_{J}\cos{\phi_{\text{DC}}}\right)\\ p_{10}=1+\beta_{p}+2\beta_{J}\cos{\phi_{\text{DC}}}\\ \Rightarrow\omega_{L}(\beta_{p},\phi_{\text{DC}})=p_{10}\;\omega_{L}(0,\phi_{\text{DC}})\end{gathered}$

### v3

@@ Appendix E Effects of parasitic loop inductance
where $l_{m}$, $c_{m}$ and $r_{m}$ are the m-th order derivatives of the potential energy functions of the left, shunting and right LINC branches, respectively, evaluated at $\pm\frac{\phi}{2}$. Finally, for driven behavior, we can also compute the n-th order derivative with respect to external flux, giving the general expansion coefficient
@@
$\left\{\begin{aligned} U_{L,R}&=\frac{E_{\text{loop}}}{2}\left(\beta_{l}\sin\varphi_{J_{L,R}}\right)^{2}-E_{J}\cos\varphi_{J_{L,R}}\\ \phi_{L,R}&=\varphi_{J_{L,R}}+\beta_{l}\sin\varphi_{J_{L,R}}\end{aligned}\right$. (51)
@@
[…]
Appendix F Effect of a series parasitic inductance
We now consider the effect of a parasitic inductance in series with the LINC loop, which may arise from the geometric inductance of the metallic arms of the coupler that form the electric dipole necessary for self-capacitance and coupling. As in appendix A, we label the phase across the LINC as $\theta_{c}$ (with the flux threading it as $\phi_{d}$), with potential energy $U_{\mathrm{LINC}}$. We then consider the phase across the whole effective dipole (LINC ++ parasitic inductor) as $\theta$, with corresponding potential energy $U_{\text{tot}}$. Each nonlinear process activated by the coupler is then described through derivatives of this total potential energy, and can be analytically derived. We will see that for a symmetric LINC, this parasitic inductance will primarily deform the potential and dilute the nonlinearity, but not shift the potential minima – thus preserving the LINC’s linearity at $\phi_{\text{DC}}=\pi/2$. We outline our derivation in a pedagogical manner below, with the actual analytic results for the LINC Hamiltonian detailed in section F.3.
@@ Appendix F Effect of a series parasitic inductance
To evaluate the correspondence in Eq. (53), we appeal to the current conservation condition $i=i_{\mathrm{LINC}}$ to relate first-order partial derivatives of $U_{\text{tot}}$ and $U_{\mathrm{LINC}}$, following methodology similar to [29, 32, 69]. As $i=\frac{2\pi}{\Phi_{0}}\frac{\partial U_{\text{tot}}}{\partial\theta}$, and $i_{\mathrm{LINC}}=\frac{2\pi}{\Phi_{0}}\frac{\partial U_{\mathrm{LINC}}}{\partial\theta_{c}}$, the following relation always holds:
@@
F.1 Representing the total potential energy and phase
In this section, we analyze the change of variables from the LINC variables $(\theta_{c},\phi_{d})$ to the effective dipole variables $(\theta,\phi_{d})$ and study its invertibility. We define the bi-variate function implementing the change of variables as $\mathbf{F}$, such that $(\theta,\phi_{d})=\mathbf{F}(\theta_{c},\phi_{d})$. Let us define the superconducting phase drop across the parasitic inductor to be $\theta_{p}$, satisfying $\theta=\theta_{c}+\theta_{p}$, and its current to be $i_{p}$. Our derivation will primarily use the current conservation relation $i_{\mathrm{LINC}}=i_{p}$.
@@ F.1 Representing the total potential energy and phase
Applying current conservation and phase addition to eqs. 55 and 56, the effective dipole phase reads,
@@
F.2 Computing derivatives of the total potential
@@
$\left\{\begin{aligned} &1=\frac{\partial\theta}{\partial\theta_{c}}p_{10}\\ &0=\frac{\partial\theta}{\partial\theta_{c}}p_{01}+\frac{\partial\theta}{\partial\phi_{d}}.\end{aligned}\right$. (63)
@@
$\left\{\begin{aligned} p_{10}&=\frac{1}{1+\beta_{p}+\beta_{J}u_{20}}\\ p_{01}&=-\frac{\beta_{J}u_{11}}{1+\beta_{p}+\beta_{J}u_{20}}=-\beta_{J}p_{10}u_{11}.\end{aligned}\right$. (65)
@@
$U_{mn}\mathrel{\mathop{\ordinarycolon}}=\frac{\partial^{m+n}U_{\text{tot}}}{\partial\theta^{m}\phi_{d}^{n}}$ (68)
@@
$U_{m+1n}=\frac{\partial^{m+n}}{\partial\theta^{m}\phi_{d}^{n}}(E_{L}p_{00}+E_{J}u_{10})$ (69)
@@
In combination with the expression for $p_{20}$ previously obtained, this gives an expression of $U_{12}$ that again only containes $p_{10}$ and $u_{mn}$ terms. Thus by recursively substituting the expressions for $p_{01}$, $p_{11}$ and $p_{20}$, any partial derivative of $U_{\text{tot}}$ can be expressed as a multinomial function of $p_{10}$ and the set of nonlinear coefficients $u_{mn}$.
@@
F.3 The LINC Hamiltonian with parasitic inductance
Given the derivatives of the total potential $U_{\text{tot}}$ with respect to $\hat{\theta}$ and $\phi_{d}$, we can find the relevant coupler Hamiltonian in the presence of linear inductance, which we derive here. Similar to eq. 21, the Hamiltonian we derive describes the equations of motion corresponding to the bosonic operator $\hat{\theta}=\theta_{zpf}\left[\phi_{\text{DC}}\right](\hat{c}+\hat{c}^{\dagger})$, which can be expanded in terms of $U_{mn}$. Note that the parasitic inductance also affects $U_{2,0}$ and hence the spread of the wavefunction $\theta_{zpf}$, whose intrinsic flux dependence affects the every flux driven process.
First, let us consider the static LINC potential, given by:
$\begin{split}U_{\text{LINC}}&=\sum_{m}\tilde{g}_{m,0}\left(\hat{c}+\hat{c}^{\dagger}\right)^{m}\\ \tilde{g}_{m,0}&=U_{m,0}\left.\frac{\theta_{zpf}^{m}}{m!}\right|_{\phi_{\text{DC}}}.\end{split}$
@@ F.3 The LINC Hamiltonian with parasitic inductance
$\begin{gathered}U_{2,0}=p_{10}\left(E_{L}+2E_{J}\cos{\phi_{\text{DC}}}\right)\\ p_{10}=\left(1+\beta_{p}+2\beta_{J}\cos{\phi_{\text{DC}}}\right)^{-1}\\ \Rightarrow\omega_{L}(\beta_{p},\phi_{\text{DC}})=\sqrt{p_{10}}\;\omega_{L}(0,\phi_{\text{DC}})\end{gathered}$